OpenStudy (jayden):

Find the value of y so that so that the line passing through (4, y) and (5, 3) has a slope of 1.

OpenStudy (jayden):

could any one help I have tried but does not make sens to me thankyou

OpenStudy (anonymous):

You already know the slope, you're just missing a y coordinate. so plug everything in and solve for that 1=(3-y)/(5-4)

OpenStudy (anonymous):

Can you do it?

OpenStudy (jayden):

thank you for your help got sidetracked.

OpenStudy (jayden):

so its 1= 15-12-5y+4y

OpenStudy (anonymous):

1=(3-y)/1 1=3-y -2=-y 2=y

OpenStudy (jayden):

y= -28

OpenStudy (jayden):

ow i see my mistake now thankyou for your help

OpenStudy (jayden):

-2=x-9 x=7 right?

OpenStudy (anonymous):

What?

OpenStudy (jayden):

The question is Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2.

OpenStudy (jayden):

-2=x-9/-2

OpenStudy (anonymous):

-2=(9-5)/(x-10) and solve for x remember the slope formula up there it refers to 2 points (x1,y1) (x2,y2)

OpenStudy (anonymous):

-2=4/(x-10) -2(x-10)=4 -2x+20=4 -2x=-16 x=8